# Feasibility of Learning Weighted Automata on a Semiring

### v3

@@ Appendix B Examples
Finally, $f_{2}$ is not guessable. Suppose for contradiction that $f_{2}$ is guessable. Then there exist finite sets $Q$, $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$. Without loss of generality, we may assume that for a fixed constant $N$ (to be chosen below) we have $n>N$ where $n$ is the length of the longest word in $T$ (since for all $T^{\prime}\supseteq T$ we also have $\Lambda_{Q,T}=\Lambda_{Q,T^{\prime}}=\Lambda_{Q}\neq\emptyset$). Hence for all $\lambda$ in $\Lambda_{Q,T}$, we have that $\mathcal{H}_{Q,T,\lambda}$ computes $f_{2}$. If $Q$ only contains words starting with an $a$, then for all $q\in Q$, $\langle qa\rangle_{T}=\langle q\rangle+1$ and $\langle qb\rangle_{T}=\langle q\rangle$, and it is easy to see that a hypothesis automaton contructed from $Q$, $T$ and a $\lambda$ of this shape would not compute $f_{2}$. A symmetric argument holds if $Q$ only contains words starting with $b$. If $Q$ contains a word $q$ starting with $a$ and a word $q^{\prime}$ starting with $b$, then, we may write $\langle qa\rangle_{T}=\max(\langle q\rangle_{T}+1,\langle q^{\prime}\rangle_{T}-n)$, $\langle qb\rangle_{T}=\max(\langle q\rangle_{T},\langle q^{\prime}\rangle_{T}-n)$, $\langle q^{\prime}a\rangle_{T}=\max(\langle q\rangle_{T}-n,\langle q^{\prime}\rangle_{T})$ and $\langle q^{\prime}b\rangle_{T}=\max(\langle q\rangle_{T}-n,\langle q^{\prime}\rangle_{T}+1)$. It is then easy to see that there exists a hypothesis automaton constructed from sets $Q$, $T$ and a $\lambda$ using the values from the linear combinations above that would not compute $f_{2}$; it would for example misbehave on a word of the form $qa^{n}bb^{2n}$. Indeed, since $q$ starts with $a$ the target function computes $|q|_{a}+n$. If there is no run from any initial state to $q$ in the hypothesis automaton, then the automaton computes $\infty$- ∞ and we are done. Otherwise, there is a run of weight $x+2n$, where $x$ is the maximal weight of a run from an initial state to $q$. Taking $N=|q|_{a}-x$ we see that $x+2n>x+n+N=|q|_{a}+n=f_{2}(qa^{n}bb^{2n})$.
@@
Since $\mathbb{Z}_{\max}$ and $\mathbb{R}_{\max}$ are commutative, and $f_{2}$ is weakly guessable but neither guessable nor weakly co-guessable, it follows that $\bar{f_{2}}$ is weakly co-guessable but neither co-guessable nor weakly guessable (since a function over a commutative semiring is (weakly) guessable if and only if its mirror is (weakly) co-guessable).
@@
For $S=\mathbb{N}_{\max}$, consider the function $f_{3}$ computed by the automaton depicted in Figure 1(a) and $F$ its Hankel matrix. It is easy to see that $f_{3}$ is computed by a literal automaton, by pushing the weights on the first transitions in the automaton depicted in the figure, as final weights. So $f_{3}$ is weakly guessable and hence strongly guessable. Let us prove now that it is not weakly co-guessable. For all words $w$, we have $f_{3}(aw)=|aw|_{a}$, and $f_{3}(bw)=|bw|_{b}$. Suppose for contradiction that $f_{3}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=|w_{i}|_{a}+1$ and $F_{b,w_{i}}=|w_{i}|_{b}+1$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=N+1$ and $F_{b,w}=1$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
@@
For $S=\mathbb{N}$, consider the function $f^{\prime}_{3}$ computed by the automaton depicted in Figure 1(c) and $F$ its Hankel matrix. It is easy to see that $f^{\prime}_{3}$ is computed by a literal automaton, by pushing the weights on the first transitions in the automaton depicted in the figure, as final weights. So $f^{\prime}_{3}$ is weakly guessable and hence strongly guessable. Let us prove now that it is not weakly co-guessable. For all words $w$, we have $f^{\prime}_{3}(aw)=2^{|aw|_{a}}$, and $f^{\prime}_{3}(bw)=2^{|bw|_{b}}$. Suppose for contradiction that $f^{\prime}_{3}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=2^{|w_{i}|_{a}+1}$ and $F_{b,w_{i}}=2^{|w_{i}|_{b}+1}$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=2^{N+1}$ and $F_{b,w}=2$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
@@
For $S=\mathcal{P}_{\rm fin}(\Sigma^{*})$, consider the function $f^{\prime\prime}_{3}$ computed by the automaton depicted in Figure 1(e) and $F$ its Hankel matrix. It is easy to see that $f^{\prime\prime}_{3}$ is computed by a literal automaton, by pushing the weights on the first transitions in the automaton depicted in the figure, as final weights. So $f^{\prime\prime}_{3}$ is weakly guessable and hence strongly guessable. Let us prove now that it is not weakly co-guessable. For all words $w$, we have $f^{\prime\prime}_{3}(aw)=a^{|aw|_{a}}$, and $f^{\prime\prime}_{3}(bw)=b^{|bw|_{b}}$. Suppose for contradiction that $f^{\prime\prime}_{3}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=a^{|w_{i}|_{a}+1}$ and $F_{b,w_{i}}=b^{|w_{i}|_{b}+1}$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=a^{N+1}$ and $F_{b,w}=b$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
@@
Finally, we give an example of a function over $S=\mathbb{R}_{\geq 0}$ that is strongly guessable but not weakly co-guessable. Consider the automaton obtained from that in Figure 1(c) by replacing all weights labelled by $b$ by $1$. In a similar manner to the previous cases, it is easy to see that the function computed, which we denote by $f_{5}$, is computed by a literal automaton and hence is weakly guessable. For all words $w$, we have $f_{5}(aw)=2^{|aw|_{a}}$, and $f_{5}(bw)=1$. Taking $Q=T=\{\varepsilon,a,b\}$ it is straightforward to verify that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$ (indeed, for each of $\langle\varepsilon\rangle,\langle qa\rangle,\langle qb\rangle$ where $q\in Q$ we find that there is exactly one way to write the given vector as a linear combination of the $\langle q^{\prime}\rangle$ with $q^{\prime}\in Q$). Hence $f_{5}$ is strongly guessable. Suppose then for contradiction that $f_{5}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\sum_{i}2^{|w_{i}|_{a}+1}$. Since for all $i$, we have $F_{a,w_{i}}=2^{|w_{i}|_{a}+1}$ and $F_{b,w_{i}}=1$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=2^{N+1}$ and $F_{b,w}=1$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
@@
Since $\mathbb{N}_{\max}$ and $\mathbb{N}$ are commutative, and $f_{3}$ and $f^{\prime}_{3}$ are strongly guessable but not weakly co-guessable, then $\bar{f_{3}}$ and $\bar{f^{\prime}_{3}}$ are strongly co-guessable but not weakly guessable (using once more the fact that a function over a commutative semiring is (weakly/strongly) guessable if and only if its mirror is (weakly/strongly) co-guessable).

### v4

@@ Appendix B Examples
Finally, $f_{2}$ is not guessable. Suppose for contradiction that $f_{2}$ is guessable. Then there exist finite sets $Q$, $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$. and hence for all $\underline{\lambda}$ in $\Lambda_{Q,T}$, we have that $\mathcal{H}_{Q,T,\lambda}$ computes $f_{2}$. By the observations at the start of this subsection, we must have that $Q$ contains $\varepsilon$ and at least one word, say $q$, beginning with $a$ and at least one word, say $q^{\prime}$, beginning with $b$. Without loss of generality, we may assume that for a fixed constant $N$ (to be chosen below) we have that the length of the longest word in $T$ is greater than $N$ (since for all $T^{\prime}\supseteq T$ we also have $\Lambda_{Q,T}=\Lambda_{Q,T^{\prime}}=\Lambda_{Q}\neq\emptyset$). Now let $n$ be equal to the maximal sum of the length a word in $Q$ and the length of a word in $T$ (which by the above observation can be made arbitrarily large, by modifying $T$ if necessary). For this choice of $n$ and for all $t\in T$ we have
$\langle qa\rangle_{t}$ $=$ $f_{2}(qat)=|qat|_{a}=|qt|_{a}+1\geq 0\geq|q^{\prime}t|_{b}-n\Rightarrow\langle qa\rangle_{T}=(1\otimes\langle q\rangle_{T})\oplus(-n\otimes\langle q^{\prime}\rangle_{T})$
$\langle qb\rangle_{t}$ $=$ $f_{2}(qbt)=|qbt|_{a}=|qt|_{a}+0\geq 0\geq|q^{\prime}t|_{b}-n\Rightarrow\langle qb\rangle_{T}=(0\otimes\langle q\rangle_{T})\oplus,(-n\otimes\langle q^{\prime}\rangle_{T})$
$\langle q^{\prime}a\rangle_{t}$ $=$ $f_{2}(q^{\prime}at)=|q^{\prime}at|_{b}=|qt|_{b}+0\geq 0\geq|qt|_{a}-n\Rightarrow\langle q^{\prime}a\rangle_{T}=(-n\otimes\langle q\rangle_{T})\oplus 0\langle q^{\prime}\rangle_{T}$
$\langle q^{\prime}b\rangle_{t}$ $=$ $f_{2}(q^{\prime}bt)=|q^{\prime}bt|_{b}=|q^{\prime}t|_{b}+1\geq 0\geq|qt|_{a}-n\Rightarrow\langle q^{\prime}b\rangle_{T}=(-n\otimes\langle q\rangle_{T})\oplus(1\otimes\langle q^{\prime}\rangle_{T}).$
Now let $\mathcal{H_{Q,T,\underline{\lambda}}}$ be the hypothesis automaton constructed from an element $\underline{\lambda}\in\Lambda_{Q,T}=\Lambda_{Q}$ using the linear combinations above (containing the negative coefficients $-n$) and let $\mathcal{H}_{q}$ denote the automaton obtained from $\mathcal{H}_{Q,T,\underline{\lambda}}$ by making $\langle q\rangle$ the unique initial state. Since $\underline{\lambda}\in\Lambda_{Q}$, it follows from the first part of the proof of Theorem 2.12 that for all words $w$, and all $q\in Q$, $\langle q\rangle_{w}$ is equal to the value computed by $\mathcal{H}_{q}$ on input $w$. Thus in particular we have that $f_{2}(qa^{n}bb^{2n})$ is equal to the value computed by $\mathcal{H}_{q}$ on input $a^{n}bb^{2n}$. Consider the run composed of: the transition from $q$ to $q$ labelled by $a$ for each instance of $a$ in the factor $a^{n}$, the transition from $q$ to $q^{\prime}$ labelled by $b$ for the first factor of $b$, and the transition from $q^{\prime}$ to $q^{\prime}$ labelled by $b$ for each instance of $b$ in the remaining factor $b^{2n}$. Since the final weight at state $q^{\prime}$ is (by definition) $\langle q^{\prime}\rangle_{\varepsilon}=f_{2}(q^{\prime})=|q^{\prime}|_{b}$, it then follows from the equations above (recalling that $\otimes$ denotes usual addition in the semirings under consideration) that this run has weight $n+(-n)+n+|q^{\prime}|_{b}=2n+|q^{\prime}|_{b}$ giving $f_{2}(qa^{n}bb^{2n})>2n$ (recalling that the addition of this semiring is maximisation). On the other hand, since $q$ starts with the letter $a$, we have $f_{2}(qa^{n}bb^{2n})=|q|_{a}+n$ (from the definition of this function). But now taking $N=|q|_{a}$ we have $f_{2}(qa^{n}bb^{2n})>2n>n+N=|q|_{a}+n=f_{2}(qa^{n}bb^{2n})$, giving a contradiction.
@@
Since $\mathbb{Z}_{\max}$ and $\mathbb{R}_{\max}$ are commutative, and $f_{2}$ is weakly guessable but neither guessable nor weakly co-guessable, it follows from Lemma 4.3 that $\bar{f_{2}}$ is weakly co-guessable but neither co-guessable nor weakly guessable (since a function over a commutative semiring is (weakly) guessable if and only if its mirror is (weakly) co-guessable).
@@
For $S=\mathbb{N}_{\max}$, consider the function $f_{3}$ computed by the automaton depicted in Figure 1(a) and $F$ its Hankel matrix. Thus $f_{3}(w)=|w|_{a}$ if $w$ starts with an $a$, $f_{3}(w)=|w|_{b}$ if $w$ starts with an $b$, and $f_{3}(\varepsilon)=-\infty$, and (by the same argument as for $f_{2}$) we find that $f_{3}$ is computed by a literal automaton and weakly guessable; in this case however Lemma 4.4 applies to give that $f_{3}$ is also strongly guessable (and note that the proof that $f_{2}$ is not guessable made use of negative coefficents that we do not have access to here!)) Let us prove now that it is not weakly co-guessable. For all words $w$, we have $f_{3}(aw)=|aw|_{a}$, and $f_{3}(bw)=|bw|_{b}$. Suppose for contradiction that $f_{3}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=|w_{i}|_{a}+1$ and $F_{b,w_{i}}=|w_{i}|_{b}+1$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=N+1$ and $F_{b,w}=1$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
@@
For $S=\mathbb{N}$, consider the function $f^{\prime}_{3}$ computed by the automaton depicted in Figure 1(c) and $F$ its Hankel matrix. Thus $f^{\prime}_{3}(w)=2^{|w|_{a}}$ if $w$ starts with an $a$, $f_{3}(w)=2^{|w|_{b}}$ if $w$ starts with an $b$, and $f_{3}(\varepsilon)=0$. It is easy to see that $f^{\prime}_{3}$ is computed by a literal automaton, by pushing the weights on the first transitions in the automaton depicted in the figure, as final weights. So $f^{\prime}_{3}$ is weakly guessable and hence strongly guessable. Let us prove now that it is not weakly co-guessable. For all words $w$, we have $f^{\prime}_{3}(aw)=2^{|aw|_{a}}$, and $f^{\prime}_{3}(bw)=2^{|bw|_{b}}$. Suppose for contradiction that $f^{\prime}_{3}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=2^{|w_{i}|_{a}+1}$ and $F_{b,w_{i}}=2^{|w_{i}|_{b}+1}$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=2^{N+1}$ and $F_{b,w}=2$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
@@
For $S=\mathcal{P}_{\rm fin}(\Sigma^{*})$, consider the function $f^{\prime\prime}_{3}$ computed by the automaton depicted in Figure 1(e) and $F$ its Hankel matrix. Thus $f^{\prime\prime}_{3}(w)=\{x^{|w|_{x}}\}$ if $w$ starts with an $x\in\{a,b\}$, and $f^{\prime\prime}_{3}(\varepsilon)=\emptyset$. It is easy to see that $f^{\prime\prime}_{3}$ is computed by a literal automaton, by pushing the weights on the first transitions in the automaton depicted in the figure, as final weights. So $f^{\prime\prime}_{3}$ is weakly guessable and hence strongly guessable. Let us prove now that it is not weakly co-guessable. For all words $w$, we have $f^{\prime\prime}_{3}(aw)=a^{|aw|_{a}}$, and $f^{\prime\prime}_{3}(bw)=b^{|bw|_{b}}$. Suppose for contradiction that $f^{\prime\prime}_{3}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=a^{|w_{i}|_{a}+1}$ and $F_{b,w_{i}}=b^{|w_{i}|_{b}+1}$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=a^{N+1}$ and $F_{b,w}=b$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
@@
Finally, we give an example of a function over $S=\mathbb{R}_{\geq 0}$ that is strongly guessable but not weakly co-guessable. Consider the automaton obtained from that in Figure 1(c) by replacing all weights labelled by $b$ by $1$. That is, the function computed, which we denote by $f_{5}$, is defined by $f_{5}(w)=2^{|w|_{a}}$ if $w$ begins with $a$, $f_{5}(w)=1$ if $w$ begins with $b$ and $f_{5}(\varepsilon)=0$. In a similar manner to the previous cases, it is easy to see that $f_{5}$ is computed by a literal automaton and hence is weakly guessable. Taking $Q=T=\{\varepsilon,a,b\}$ it is straightforward to verify that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$ (indeed, for each of $\langle\varepsilon\rangle,\langle qa\rangle,\langle qb\rangle$ where $q\in Q$ we find that there is exactly one way to write the given vector as a linear combination of the $\langle q^{\prime}\rangle$ with $q^{\prime}\in Q$). Hence $f_{5}$ is strongly guessable. Suppose then for contradiction that $f_{5}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\sum_{i}2^{|w_{i}|_{a}+1}$. Since for all $i$, we have $F_{a,w_{i}}=2^{|w_{i}|_{a}+1}$ and $F_{b,w_{i}}=1$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=2^{N+1}$ and $F_{b,w}=1$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
@@
Since $\mathbb{N}_{\max}$ and $\mathbb{N}$ are commutative, and $f_{3}$ and $f^{\prime}_{3}$ are strongly guessable but not weakly co-guessable, then (by Lemma 4.3) $\bar{f_{3}}$ and $\bar{f^{\prime}_{3}}$ are strongly co-guessable but not weakly guessable .